OpenStudy (alibaby):

The distance of planet Mercury from the Sun is approximately 5.8 ⋅ 10^7 kilometers, and the distance of Earth from the Sun is 1.5 ⋅ 10^8 kilometers. About how many more kilometers is the distance of Earth from the Sun than the distance of Mercury from the Sun? 4.3 ⋅ 10^7 kilometers 9.2 ⋅ 10^7 kilometers 9.2 ⋅ 10^8 kilometers 5.7 ⋅ 10^9 kilometers

OpenStudy (alibaby):

@rishavraj sorry i keep tagging you it's just your really smart ^

OpenStudy (alibaby):

it's A pr B it's one of the 2

OpenStudy (johnweldon1993):

mmhmm, one of those! \[\large 1.5\times 10^8 - 5.8\times 10^7 = 9.2\times 10^7\]